OpenStudy (anonymous):

How to solve this equation? 6^(3x-2) = 8*6^(2x)

OpenStudy (zzr0ck3r):

hint: 8 = (6^(3x-2))/6^(2x) 8 = 6^(3x-2-2x) 8 = 6^(x-2) can you do the rest?

OpenStudy (zzr0ck3r):

if a = b^c then ln(a) = ln(b^c) = c*ln(b)

OpenStudy (zzr0ck3r):

@donmenol this making sense?

OpenStudy (anonymous):

hmmm yes